# Stable Quantum Vortices in Lee-Huang-Yang Dipolar Superfluids

### v1

@@ III.4 The effect of the rotational frequency
Table 1: The effect of rotation frequency $\Omega$ on chemical potential ($\mu$), total energy per particle ($E$), number of visible vortices ($N_{v}$), and angular momentum per particle, $\left\langle L_{z}\right\rangle$ for five different sets of parameters considered in this work. Case I (the pure MF contact interaction, with $g=100$); case II (the pure DDI, with $g_{\mathrm{dd}}=100$); case III (with the contact-MF and LHY terms, without DDI, for $g=100$ and $\eta=200$); case IV (the nonzero contact-MF, DDI, and LHY terms, with $g=100$, $g_{\mathrm{dd}}=100$, and $\eta=200$); and case V (the pure LHY nonlinearity, with $\eta=200)$. In the pure-LHY superfluid, one should observe that the critical frequencies $\Omega_{c}$ for producing one and two vortex states are about the same within two decimal digits (being, respectively, $0.6401$ and $0.6410$). It implies that a single-vortex production is numerically quite challenging. In this table, the corresponding values for $N_{v}=1$ are: $\Omega=0.6401$, $\mu=1.78$, $E=1.21$, and $\langle L_{z}\rangle=1$.
@@ I II III IV V I II III IV V I II III IV V I II III IV V
The findings for the vortex configurations are summarized in Table 1 for all five nonlinearity cases that we have considered. In the table, we report the numerical results obtained for the chemical potential, total energy per particle $E$, and number of visible vortices $N_{v}$, for specific values of the rotation frequency $\Omega$, which are given with one decimal digit. For the five cases being considered, the corresponding behaviors of $E$ and $\mu$, as functions of $\Omega$, are presented in two panels of Fig. 10. The energy cost of generating vortices can be verified from the results provided in Table 1. In particular, by focusing on case V of the table, where we have only the LHY nonlinear term, the energy cost of producing a pair of vortices, as compared with the no-rotating case ($E(\Omega=0)$), is $\approx 22\%$ at $\Omega=0.7$. On the other hand, when we have only the MF contact nonlinearity (column I in the table), the cost is $\approx 14\%$ at $\Omega=0.6$. This indicates that the vortices are more likely to be produced in the presence of the pure MF contact nonlinearity than in the case of the pure LHY nonlinearity.
@@
We conclude this section by showing that all the obtained results can be well fitted by the Feynman prediction of the linear behavior of the angular momentum per particle versus the number of vortices [13], with the results provided in Fig. 11.
@@ IV Conclusion
In the present work, considering the rotating dipolar BEC, we investigate the impact of quantum fluctuations, represented by the LHY correction to the GP equation, on the formation and dynamics of individual vortices and vortex lattices. The main outcome of our analysis is related to the scenario of vortex production when the only source of nonlinearity is provided by the LHY term. This condition may approximately hold for the single-component dipolar BEC, with the MF (mean-field) contact interaction and DDI (dipole-dipole interaction) roughly canceling each other. To achieve this setting, one has to select appropriate values of the strengths of the MF contact and dipole-dipole interactions, which can be provided through the Feshbach resonance, and by adjusting the angle that determines the orientation of the atomic magnetic moments. It is shown that the LHY-only nonlinearity can be made strong enough to produce vortices in the dipolar superfluid. For the smaller numbers of stable vortices ($N_{v}\leq 5$), it being more likely to obtain even rather than odd numbers of vortices when varying the rotational frequency $\Omega$. This means that the range of $\Omega$ is relatively large to generate even numbers ($2$ or $4$) of the vortices, being much smaller for the odd numbers ($1$ and $3$). This feature is particularly pronounced in the pure LHY case, so that, for the single-vortex case, the corresponding range of the rotation frequency $\Omega$ is so narrow that it can be identified with the critical one $\Omega_{c}$ for the production of the single vortex. This effect, related to the quartic LHY nonlinearity, can also be partially attributed to the aspect ratio of the 3D trap. In a less pronounced form, a similar effect was also noted for the relatively small values of vortices in the case of pure MF contact interactions. A more detailed investigation related to this effect, considering different parametrization, may be required to clarify the role of the aspect ratio and its interplay with the nonlinearity.
@@
We have also computed, within the framework of our quasi-2D model, the number of vortices, chemical potential, and energy per particle, expressed as functions of the rotation frequency $\Omega$. For that purpose, we have considered specific combinations of the strengths of the MF contact interaction, DDI, and LHY nonlinearity. These parameters are assumed to be large enough to study the vortex production from the perspective of possible experimental realizations in dipolar BEC systems. The results demonstrate that, while the vortex numbers consistently increase with the growth of $\Omega$, the chemical potential and energy decrease. From this study, when considering the pure LHY nonlinearity, exemplified by a particular value of the strength, which can be further generalized, we conclude how to precisely identify the single-vortex critical frequency, along with the corresponding values of the chemical potential and energy. In all the investigated cases, we have observed results consistent with the Feynman’s prediction of linear dependence between angular momentum per particle and the number of vortices.

### v2

@@ III.4 The effect of the rotational frequency
Table 1: The effect of rotation frequency $\Omega$ on chemical potential ($\mu$), total energy per particle ($E$), number of visible vortices ($N_{v}$), and angular momentum per particle, $\left\langle L_{z}\right\rangle$ for five different sets of parameters considered in this work. Case I (the pure MF contact interaction, with $g=100$); case II (the pure DDI, with $g_{dd}=100$); case III (with the contact-MF and LHY terms, without DDI, for $g=100$ and $\eta=200$); case IV (the nonzero contact-MF, DDI, and LHY terms, with $g=100$, $g_{dd}=100$, and $\eta=200$); and case V (the pure LHY nonlinearity, with $\eta=200)$. In the pure-LHY superfluid, one should observe that the critical frequencies $\Omega_{c}$ for producing one and two vortex states are about the same within two decimal digits (being, respectively, $0.6401$ and $0.6410$). It implies that a single-vortex production is numerically quite challenging. In this table, the corresponding values for $N_{v}=1$ are: $\Omega=0.6401$, $\mu=1.78$, $E=1.21$, and $\langle L_{z}\rangle=1$.
@@ I II III IV V I II III IV V I II III IV V I II III IV V
The findings for the vortex configurations are summarized in Table 1 for all five nonlinearity cases that we have considered. In the table, we report the numerical results obtained for the chemical potential, total energy per particle $E$, and number of visible vortices $N_{v}$, for specific values of the rotation frequency $\Omega$, which are given with one decimal digit. For the five cases being considered, the corresponding behaviors of $E$ and $\mu$, as functions of $\Omega$, are presented in two panels of Fig. 10. The energy cost of generating vortices can be verified from the results provided in Table 1. In particular, by focusing on case V of the table, where we have only the LHY nonlinear term, the energy cost of producing a pair of vortices, as compared with the non-rotating case ($E_{\Omega=0}$), is $\approx 22\%$ at $\Omega=0.7$. On the other hand, when we have only the MF contact nonlinearity (column I in the table), the cost is $\approx 14\%$ at $\Omega=0.6$. This indicates that the vortices are more likely to be produced in the presence of the pure MF contact nonlinearity than in the case of the pure LHY nonlinearity.
@@
Figure 11 summarizes the relationship between the angular momentum per particle and the number of vortices $N_{v}$. The data exhibit a clear linear trend, in excellent agreement with the Feynman’s prediction for the angular momentum in a rapidly rotating superfluid [13]. The vortex core (where the density is zero) is a nodal line of the wave function. Preventing the kinetic energy from diverging, the core provides a cutoff, with a radius on the order of the healing length $\xi=\hbar/\sqrt{2mng_{eff}}$, where $g_{eff}$ is the effective interaction strength. Next, we briefly discuss how the different kind of nonlinear interactions affect the corresponding healing lengths.
III.5 The healing length in the LHY-corrected case
From the density results presented in Figs. 2 and 6-7, by comparing cases with identical numbers of vortices, one can verify that the vortex healing lengths (radial sizes) slightly increase when proceeding from the usual cubic (pure contact) to the pure quartic LHY nonlinearity in the GP formalism. We can estimate the healing lengths more precisely using the results in Table 1. As shown in Appendix A, the full-dimensional healing length for the LHY-corrected case is given by $\xi={\hbar}/{\sqrt{2m\left(g_{0}n_{0}+\frac{3}{2}g_{LHY}n_{0}^{3/2}\right)}}$, where $n_{0}$ is the equilibrium bulk density. So, for the usual MF, we have $\xi_{MF}={\hbar}/{\sqrt{2m\left(g_{0}n_{0}\right)}}={\hbar}/{\sqrt{2m\mu_{MF}}}$; whereas, for pure LHY interaction, $\xi_{LHY}={\hbar}/{\sqrt{3m\;g_{LHY}n_{o}^{3/2}}}={\hbar}/{\sqrt{3m\mu_{LHY}}}$. Using the dimensionless chemical potentials from Table 1, we next compare the four-vortex configurations, where the vorticity is the most stable. For the pure contact interaction: $\Omega=0.7$ (Fig. 2c), with $\mu_{MF}\approx 2.24$; and, for the pure LHY, $\Omega=0.8$ (Fig. 6c), with $\mu_{LHY}\approx 1.23$. Therefore, $\xi_{LHY}/\xi_{MF}\approx\sqrt{2\mu_{MF}/(3\mu_{LHY})}\approx 1.10$, which indicates that the vortices in pure LHY quantum fluids have healing lengths about 10% larger than for the cases with only contact interactions. With the healing length slightly larger in the LHY fluid, the space between the radial borders of the vortices will be reduced accordingly. The intervortex distance (between the central cores), as shown in Ref. [105], is given by $l_{\Omega}=\sqrt{\hbar/(m\Omega)}$, which depends on the rotational frequency, consistent with the Feynman’s prediction [13]. Therefore, when comparing cases having the same number of vortices but different nonlinearities, this intervortex distance should be slightly changed according to the changes of the rotational frequency. In our example of four vortices, the estimated distance between vortices is $l_{\Omega_{LHY}}/l_{\Omega_{MF}}\approx\sqrt{7/8}\approx 0.935$.
@@ IV Conclusion
In the present work, considering the rotating dipolar BEC, we investigate the impact of quantum fluctuations, represented by the LHY correction to the GP equation, on the formation and dynamics of individual vortices and vortex lattices. The main outcome of our analysis is related to the scenario of vortex production when the only source of nonlinearity is provided by the LHY term. This condition may approximately hold for the single-component dipolar BEC, with the mean-field (MF) contact interaction and DDI (dipole-dipole interaction) roughly canceling each other. To achieve this setting, one has to select appropriate values of the strengths of the MF contact and dipole-dipole interactions, which can be provided through the Feshbach resonance, and by adjusting the angle that determines the orientation of the atomic magnetic moments. It is shown that the LHY-only nonlinearity can be made strong enough to produce vortices in the dipolar superfluid. For the smaller numbers of stable vortices ($N_{v}\leq 5$), it being more likely to obtain even rather than odd numbers of vortices when varying the rotational frequency $\Omega$. This means that the range of $\Omega$ is relatively large to generate even numbers ($2$ or $4$) of the vortices, being much smaller for the odd numbers ($1$ and $3$). This feature is particularly pronounced in the pure LHY case, so that, for the single-vortex case, the corresponding range of the rotation frequency $\Omega$ is so narrow that it can be identified with the critical one $\Omega_{c}$ for the production of the single vortex. This effect, related to the quartic LHY nonlinearity, can also be partially attributed to the aspect ratio of the 3D trap. In a less pronounced form, a similar effect was also noted for the relatively small values of vortices in the case of pure MF contact interactions. A more detailed investigation related to this effect, considering different parametrization, may be required to clarify the role of the aspect ratio and its interplay with the nonlinearity.
@@
We have also computed, within the framework of our quasi-2D model, the number of vortices, chemical potential, and energy per particle, expressed as functions of the rotation frequency $\Omega$. For that purpose, we have considered specific combinations of the strengths of the MF contact interaction, DDI, and LHY nonlinearity. These parameters are assumed to be large enough to study the vortex production from the perspective of possible experimental realizations in dipolar BEC systems. The results demonstrate that, while the vortex numbers consistently increase with the growth of $\Omega$, the chemical potential and energy decrease. From this study, when considering the pure LHY nonlinearity, exemplified by a particular value of the strength, which can be further generalized, we conclude how to precisely identify the single-vortex critical frequency, along with the corresponding values of the chemical potential and energy. In all the investigated cases, we have observed results consistent with Feynman’s prediction of a linear dependence between angular momentum per particle and the number of vortices. Finally, we briefly discuss the vortex distribution and corresponding healing lengths in a trap in view of the different kinds of nonlinearities, by confronting the case with pure MF, with that one with pure LHY quantum fluid, with rotations such that the same number of vortices is produced. The results are consistent with experimental observations that $\xi_{LHY}$ $\xi_{MF}$.
@@ Acknowledgements.
Appendix: The healing length of a vortex in an LHY-corrected superfluid
The healing length for a vortex in an LHY-corrected superfluid is defined by the linearized density recovery around the bulk equilibrium, governed by an effective coupling $g_{eff}$, that includes the derivative of the LHY term with respect to density. In view of our purpose, to simplify this appendix, we consider a non-dipolar system. The density is zero at the core of the vortex and heals to the equilibrium density $n_{0}\equiv|\psi_{0}({\bf r})|^{2}$ as going away from it. This equilibrium density itself is determined by balancing the MF (with interaction $g_{0}$) and LHY (with interaction strength $g_{LHY}$) pressures in the bulk, often leading to a flat-top droplet if the confinement is weak. By ignoring the external potential, considering uniform gas bulk equilibrium condition, the pressure balance gives
$\mu_{0}=g_{0}n_{0}+g_{LHY}{n_{0}}^{3/2}.$ (A1)
For a stationary straight vortex line with winding number one, using the ansatz $\psi({\bf r})=\sqrt{n(r)}{\rm e}^{{\rm i}\varphi}$, with $r$ being the distance from the core, the real part of the extended GP equation is
$\hskip-5.69046pt\frac{-\hbar^{2}}{2m}\left[\frac{1}{r\sqrt{n}}\frac{d}{dr}\left(r\frac{d\sqrt{n}}{dr}\right)-\frac{1}{r^{2}}\right]+g_{0}n+g_{LHY}{n}^{3/2}=\mu_{0}.$ (A2)
At $r\to\infty$, $n\to n_{0}$, with small density deviations being $\delta n=n_{0}-n$, the linearization yields the healing length. The effective nonlinearity is now $g_{0}n+g_{LHY}n^{3/2}$. By expanding the chemical potential about $n_{0}$, we have
$\mu(n)\approx\mu_{0}+\left(g_{0}+\frac{3}{2}g_{LHY}n_{0}^{1/2}\right)(n-n_{0}),$ (A3)
such that the effective coupling strength governing the stiffness against density variations is
$g_{eff}=g_{0}+\frac{3}{2}g_{LHY}n_{0}^{1/2}.$ (A4)
For a self-bound droplet, where $g_{0}$ can be negative, but $g_{eff}$ is positive and nonsmall, the LHY term provides the dominant repulsion stabilizing the core. In analogy with the standard formalism, the healing length of a vortex in an LHY-corrected superfluid is given by
$\xi=\frac{\hbar}{\sqrt{2mg_{eff}n_{0}}}=\frac{\hbar}{\sqrt{{2m}\left(g_{0}+\frac{3}{2}g_{LHY}n_{0}^{1/2}\right)n_{0}}}.$ (A5)
Therefore, in a regime dominated by the LHY term (the droplet regime), assuming $g_{0}=0$, we obtain the healing length, as
$\xi_{LHY}\approx\frac{\hbar}{\sqrt{3mg_{LHY}n_{0}^{3/2}}}\;\;\;=\;\;\;\frac{\hbar}{\sqrt{3m\mu_{LHY}}},$ (A6)
which scales with the bulk density as $n_{0}^{-3/4}$, different from the MF scaling $\sim n_{0}^{-1/2}$. In terms of the respective chemical potentials, the factor $1/\sqrt{2}$ is replaced by $1/\sqrt{3}$.